OpenStudy (anonymous):

a pyramid has a slant height of 3 ft and square base that measures 4 ft on a slide find the Surface Area

OpenStudy (anonymous):

the sa formula of a pyramid is SA= s^2+2*s*l were s=base side length and l= length of the sides

OpenStudy (anonymous):

l = slant height

OpenStudy (anonymous):

so SA= 4^2+2*4*3

OpenStudy (anonymous):

lets simplify SA= 16+24

OpenStudy (anonymous):

SA=40

OpenStudy (anonymous):

hope this helps!!

OpenStudy (anonymous):

Thanks! :D

OpenStudy (anonymous):

np!!:)